OpenStudy (anonymous):

solve the system by teh addition method x+1=1 x-y=-5

OpenStudy (anonymous):

Subtract the second equation from the first one. left side is \[x + 1 - x + y\] right side is \[1 - (-5)\] so \[y+1 = 6\] y = 5

OpenStudy (anonymous):

You mean the subtraction method right? @m_alex75

OpenStudy (radar):

Do as Matt.Mawson suggested but multiply the 2nd equation (x-y=-5) by -1 getting -x+y= 5 now add the first equation (instead of subtract) x + 1 = 1 -x + y = 5 Adding -------------- y + 1 +6 y = 5 However this required an extra step to comply with instructions of the problem. Matt.Mawson used the most logical approach, even if it was contrary to the instructions submitted with the problem.